# ClearDepth: Enhanced Stereo Perception of Transparent Objects for Robotic Manipulation

### v1

@@ IV-B3 Evaluation on Our Transparent Object Dataset
To validate the efficacy of our model and our proposed dataset for the task of depth recovery of transparent objects using stereo vision, we fine-tuned our pre-trained model on the SynClearDepth dataset. This fine-tuning process involved utilizing the same training parameters established during the pre-training phase. Additionally, we finetuned the baseline model, RAFT-Stereo [25] and other SOTA models on Middlebury benchmark, on the SynClearDepth dataset and conducted a performance comparison. This comparative analysis aimed to underscore the advancements and improvements our model offers in the specific context of stereo-based depth perception for transparent objects. Tab. II displays the quantitative validation results. The visualization results of the stereo imaging for transparent objects are illustrated in Fig. 7. From the visualization results, it is evident that the imaging performance of our network, specifically designed for transparent objects, significantly surpasses that of other stereo imaging methods in the transparent object regions.
IV-B4 Ablation Study of Feature Post-Fusion Module
To ascertain the impact of our feature post-fusion module, we embarked on a series of ablation studies. These studies compared networks equipped with and without the feature post-fusion module, specifically focusing on our transparent object dataset, SynClearDepth. The results in Tab. III demonstrate that the inclusion of the feature post-fusion module significantly improves the network’s overall performance in processing transparent objects. This enhancement is particularly evident in challenging scenarios involving complex transparency and light refraction, underscoring the module’s effectiveness in capturing and integrating crucial features for more accurate depth estimation and object recognition in such contexts. The ablation study utilizes 100,000 steps in training.
IV-C Robotic Grasping Experiments
We employ the ZED 2 stereo camera with our proposed network, integrated with a robotic arm and a five-fingered dexterous hand, to perform grasping tasks. We tested the success rate of grasping transparent objects in service robot scenarios, following the experimental setup of FFHClutteredGrasping [47]. Using our stereo method, we generated point clouds of scenes with transparent objects and performed pose estimation. We then planned the grasping posture for the five-fingered hand. We conducted 150 experiments for each scenario, varying the number of transparent objects from 1 to 5, and the results are shown in Table IV. We achieved an average success rate of 86.2%. Our approach integrates seamlessly into existing robotic grasping frameworks without needing additional processing of intermediate results. Our method does not require multi-view images capture, thus maintaining the robot’s operational efficiency. Additionally, our data generation process is fast and efficient, enabling rapid expansion of datasets for further tasks. Our findings demonstrate that our transparent object depth recovery network significantly improves the robot’s ability to grasp transparent objects.
@@ IV-C Robotic Grasping Experiments
V Conclusion
In this paper, we propose a new stereo transparent object dataset SynClearDepth, containing RGB, depth, mask, object poses, and camera poses. Compared to other real datasets, our simulated data does not introduce the inherent noise of real 3D sensors. Unlike other simulated datasets, we have collected a large number of indoor scene models to fill the dataset background, avoiding ambiguities for depth recovery networks. We also propose a stereo depth recovery network for transparent objects, which enhances the underlying structural information, allowing our network to recover the depth of transparent objects implicitly without the need for a transparent object mask prior. We conducted comprehensive comparative experiments on our model using both public datasets and our proprietary dataset for transparent objects. Additionally, we performed ablation studies to evaluate the design of our model. These experiments demonstrated the exceptional performance of our model, highlighting its robustness and accuracy in various scenarios for transparent object perception, aiding in the manipulation tasks of transparent objects by robots.

### v2

@@ IV-C3 Evaluation on Transparent Object Dataset
To validate our model and dataset for transparent object depth recovery in stereo vision, we fine-tuned our pre-trained model on the SynClearDepth dataset using the same training parameters as pre-training. We also fine-tuned RAFT-Stereo [25] and other SOTA models from the Middlebury benchmark on SynClearDepth for comparison. This analysis highlights our model’s improvements in stereo-based depth perception for transparent objects. Tab. II presents quantitative results, while Fig. 7 visualizes stereo imaging performance. AvgErr (Average Error) and RMS (Root Mean Square Error) measure numerical error, while Bad 0.5 (%), Bad 1.0 (%), Bad 2.0 (%), Bad 4.0 (%) reflect relative error. Results show our model achieves strong performance in numerical error and outperforms all others in relative error. Our model is more efficient than others, achieving comparable performance without the high computational cost of cross-attention or multi-model ensembles. This is due to innovations in the image encoder, making our approach more suitable for robotics. Additionally, our TensorRT implementation enables real-time inference at 50 FPS on consumer GPUs, whereas other models, due to their complex designs, are impractical for deployment. We perform qualitative analysis on real-world images. Results in supplementary video show that leveraging a physically realistic renderer enables strong generalization in real world, with performance consistent across domain-shifted test sets. In this work, we adopt a stereo-based approach instead of the Realsense-based imaging methods [33]. Realsense cameras, as consumer-grade devices, do not meet industrial standards. Moreover, their open-source datasets are limited in both diversity and scale compared to stereo datasets, restricting future extensions. Additionally, Realsense IR projection penetrates transparent objects, leading to visual information loss, which we avoid to ensure robustness.
@@
IV-C4 Ablation Study of Feature Post-Fusion Module
@@
To evaluate the impact of our feature post-fusion module, we conducted ablation studies on the SynClearDepth dataset. We compared networks with and without the module, as shown in Tab. III. Results indicate a substantial performance boost, especially in handling complex transparency and light refraction, highlighting its effectiveness in enhancing depth estimation and object recognition. Each study was trained for 100,000 steps.
IV-D Robotic Grasping Experiments
We employ the ZED 2 stereo camera with our network, integrated with a robotic arm and five-fingered hand, for transparent object grasping. Following grasping pipeline from  [47], we conducted 150 trials per scenario, varying object counts from 1 to 5, achieving an 86.2% average success rate (Tab. IV). Our method seamlessly integrates into existing grasping frameworks without extra intermediate processing or multi-view capture, ensuring efficiency. Additionally, our fast data generation enables rapid dataset expansion. Results show that our depth recovery network substantially enhances transparent object grasping.
V Conclusion and Future Work
In this work, we present a complete visual perception framework for transparent object manipulation in service robotics scenarios, spanning synthetic data generation, stereo depth estimation, and real-world robotic validation. We propose an efficient real-time stereo depth recovery network that combines a cascaded vision transformer backbone with a structural feature post-fusion module, enabling fine-grained structural perception and accurate depth recovery of transparent objects without relying on mask priors. To address the data scarcity challenge in transparent object perception, we construct SynClearDepth, a high-quality simulation dataset containing diverse household environments and realistic object placements. It provides accurate RGB, depth maps, instance masks, and pose annotations, substantially enhancing model generalization in real-world scenarios. We validate our model through extensive comparisons on public and proprietary datasets, along with ablation studies. Experimental results demonstrate that our approach outperforms existing methods on both public and proprietary benchmarks, particularly in structure-aware and boundary-level depth estimation. Results demonstrate its robustness, accuracy and efficiency, supporting transparent object manipulation in robotics. Furthermore, real-world robotic grasping experiments show that our method can be seamlessly integrated into grasping pipelines without requiring multi-view capture or additional pre-processing, and achieves stable and precise manipulation of transparent objects. These results highlight the practicality and applicability of stereo-based transparent object depth estimation in real-world robotic tasks.
Expanding stereo datasets across domains is an ongoing effort. Recent research on visual foundation models highlights the critical role of high-quality synthetic data, yet the open-source community still lacks sufficient quantity and diversity. Future work will focus on expending stereo data for robotic grasping, enabling commercial deployment in logistics sorting for transparent and reflective objects like cosmetics.